Create data columns for:

a. [O2(aq)] in Molarity units
b. KH (using Henry’s Law and PO2 of 0.21 Atm)
c. Ln(KH)
d. 1/T (be sure to convert temperatures to Kelvin)

Conduct a linear regression of your ln (K) vs 1/T data; use these results to find:

a. Delta H in units of J/mole
b. Delta S in units of J/mole-K
c. Comment on the magnitude of your DH compared to the magnitude of various types of
intermolecular forces; is your answer reasonable? What specific intermolecular
interactions are responsible for this release of energy?

Temp.(deg C) O2 Sol (ppm)
0 14.62
1 14.22
2 13.83
3 13.46
4 13.11
5 12.77
6 12.45
7 12.14
8 11.84
9 11.56
10 11.29
11 11.03
12 10.78
13 10.54
14 10.31
15 10.08
16 9.87
17 9.66
18 9.47
19 9.28
20 9.09
21 8.92
22 8.74
23 8.58
24 8.42
25 8.26
26 8.11
27 7.97
28 7.83
29 7.69
30 7.56
31 7.43
32 7.3
33 7.18
34 7.06
35 6.95
36 6.84
37 6.73
38 6.62
39 6.52
40 6.41
Answer #1

a. ppm means parts per million.

1 ppm of O2 = 1 mg of O2 in 1L

No of moles = Given Wt / Mol Wt

Thus,

Moles = 0.001 g/ 32 g/mol = 3.125*10-5 moles

Thus 1 ppm = 3.125*10-5 moles/L

Using this relation, concentration of O2 for the entire data will calculated.

For instance-

14.62 ppm =  14.62* 3.125*10-5 moles/L = 4.57*10-4 M

b. As per Henry's law-

Solubility = Partial Pressure * Henry's Constant

For solubility = 4.568*10-4 M

Partial Pressure = 0.21 atm

KH = 4.568*10-4 M /0.21 atm = 2.17*10-3 M/atm

Similarly the value of KH can be determined for the entire data.

Temp (deg C) O2 (ppm) O2 (M) KH ln KH Temp (K) 1/Temp (1/K)
0 14.62 0.000457 0.002176 -6.13045 273.15 0.003660992
1 14.22 0.000444 0.002116 -6.15819 274.15 0.003647638
2 13.83 0.000432 0.002058 -6.186 275.15 0.003634381
3 13.46 0.000421 0.002003 -6.21312 276.15 0.00362122
4 13.11 0.00041 0.001951 -6.23947 277.15 0.003608154
5 12.77 0.000399 0.0019 -6.26574 278.15 0.003595182
6 12.45 0.000389 0.001853 -6.29112 279.15 0.003582303
7 12.14 0.000379 0.001807 -6.31634 280.15 0.003569516
8 11.84 0.00037 0.001762 -6.34136 281.15 0.00355682
9 11.56 0.000361 0.00172 -6.36529 282.15 0.003544214
10 11.29 0.000353 0.00168 -6.38893 283.15 0.003531697
11 11.03 0.000345 0.001641 -6.41222 284.15 0.003519268
12 10.78 0.000337 0.001604 -6.43515 285.15 0.003506926
13 10.54 0.000329 0.001568 -6.45767 286.15 0.003494671
14 10.31 0.000322 0.001534 -6.47973 287.15 0.0034825
15 10.08 0.000315 0.0015 -6.50229 288.15 0.003470415
16 9.87 0.000308 0.001469 -6.52334 289.15 0.003458413
17 9.66 0.000302 0.001438 -6.54485 290.15 0.003446493
18 9.47 0.000296 0.001409 -6.56471 291.15 0.003434656
19 9.28 0.00029 0.001381 -6.58498 292.15 0.003422899
20 9.09 0.000284 0.001353 -6.60567 293.15 0.003411223
21 8.92 0.000279 0.001327 -6.62455 294.15 0.003399626
22 8.74 0.000273 0.001301 -6.64493 295.15 0.003388108
23 8.58 0.000268 0.001277 -6.66341 296.15 0.003376667
24 8.42 0.000263 0.001253 -6.68223 297.15 0.003365304
25 8.26 0.000258 0.001229 -6.70142 298.15 0.003354016
26 8.11 0.000253 0.001207 -6.71975 299.15 0.003342805
27 7.97 0.000249 0.001186 -6.73716 300.15 0.003331667
28 7.83 0.000245 0.001165 -6.75488 301.15 0.003320604
29 7.69 0.00024 0.001144 -6.77292 302.15 0.003309614
30 7.56 0.000236 0.001125 -6.78997 303.15 0.003298697
31 7.43 0.000232 0.001106 -6.80732 304.15 0.003287851
32 7.3 0.000228 0.001086 -6.82497 305.15 0.003277077
33 7.18 0.000224 0.001068 -6.84154 306.15 0.003266373
34 7.06 0.000221 0.001051 -6.8584 307.15 0.003255738
35 6.95 0.000217 0.001034 -6.8741 308.15 0.003245173
36 6.84 0.000214 0.001018 -6.89006 309.15 0.003234676
37 6.73 0.00021 0.001001 -6.90627 310.15 0.003224246
38 6.62 0.000207 0.000985 -6.92275 311.15 0.003213884
39 6.52 0.000204 0.00097 -6.93797 312.15 0.003203588
40 6.41 0.0002 0.000954 -6.95498 313.15 0.003193358

1/T 0.0031 0.0032 0.0033 0.0034 0.0035 0.0036 0.0037 y = 1745.2x - 12.546 R = 0.9978 in Ink 1000

This is the required plot.